Input an expression for the frequency of waves
1
Expert's answer
2019-07-19T13:21:36-0400

If we consider waves on water, the frequency shows how many times the crests will be replaced with troughs and vice versa within a second if we consider a single wave.



The wavelength is "\lambda", its speed is "v", thus the period - time required for the wave to turn upside down (the wave does not always do it, but if we consider how it looks initially and finally when propagates, it seems like the wave really turned upside down):


"T=\frac{\lambda}{v},"

and the frequency is the reciprocal of the period:


"f=\frac{1}{T}=\frac{v}{\lambda}."

The frequency is measured in Hertz (Hz)